NEXT
Make New Entry,
Make Followup Entry
User name Sarah
Log entry time 18:22:50 on September 16, 2009
Entry number 289983
keyword=Run Coordinator's Report for September 15 Swing to September 16 Day
Run Coordinator's Report for September 15 Swing to September 16 Day
Summary
No beam throughout the night, and the troubles with iocha14/iocha16 and
with spot++ were fixed in the morning, so we have both arms again. We
had a few hours of beam in the morning, but it wasn't very good, and MCC
took the beam away to work out the beam loss issues in the machine for 3
- 6 hours at 2:00pm.
Significant News
- Last night's swing shift started with no beam, and with only
the left arm due to the problems with iocha16. Bob Michaels did some
work on the scalers (see halog
entry 289817 and halog
entry 289846).
- There was trouble with spot++ last night, and an access had to made
to check the DAQ cables to see if they were loose. After working in the
hut, people should always check that it is working, since it doesn't if
any cables are joggled. See halog
entry 289819 , halog
entry 289857, halog
entry 289872, and halog
entry 289894. In this case we were having trouble with both, but as
a reminder, we have spot++ and spot_R, so if one doesn't work, go ahead
and try the other.
- The iocha14/iocha16 issues were resolved! See halog
entry 289974.
- We did get beam for a short time from 9:40am to about 2:00pm, but
the quality wasn't good, and MCC took the beam away (for 3 to 6 hours) to
work on the problems (namely the large beam loss in the machine). Beam
isn't expected back until well into the Swing shift.
- Q2 quenched at 3:35pm, and was reset remotely.
- The Compton is still having issues, but we are now taking Compton
data, and the experts are working on analyzing and understanding these
data so that we have a better understanding of what is going on.
Plans
- The detailed run plan for the night is in halog
289981.
- Establish good beam (do harps, spot++, check raster), then go to
production running on the left arm.
- Take lots of production data! Watch the lumi widths to see if
anything changes. Please check the histograms on the online data monitor
with the reference ones on the shift how-to page (http://hallaweb.jlab.org/experiment/HAPPEX/opmanual/opmanual.html#HPXStartingonlineanalysis).
- Run spot++ every couple of hours and log it. Change the prescale
factor PS8 to 1 and PS1 and 2 to 99999.
- Take Compton data, starting a new run every two hours. Call the
Compton experts if the background rates get so high that they trip off
the Compton detector HV.
- We are planning to have a Moller measurement tomorrow.
- When we go back to production and things are somewhat stable, if our
lumi widths are up into the 1300-1500 range, go down to 95 or 90 muA and
take data for at least 30 minutes so that we can study the target
boiling. Don't go below 100 muA unless you see the lumi widths go up to
this range at 100 muA.
Concerns
- Ed reminded us that the tech on-call tonight is Todd Ewing.
- Scot Spiegel would like to make an access if we are going to be
down for a while at a time when we can go to controlled access (in the
daytime, 8am-5pm) to get some poly for shielding. If such a time
presents itself, his extension is x5900, his cell is 757-536-5900, and
his pager is 757-584-5900.
- If anyone must reset a magnet, they must make a halog entry
stating which magnet on which arm, and what was wrong with it to require
a reset. The word "tech" must appear in the subject line.
- Be sure that the prompt analysis is done for every run!
- Is bpm12 still saturating?
- Hall C charge asymmetry feedback.
- The BTA program doesn't subtract any time to speak of for
beam trips. Please manually subtract a full minute for each beam trip in
the beam time accounting.
- Please write a detailed shift summary (including runs taken and any
activities) and submit the BTA for each shift.
- Please make sure that the shift checklist is done once each shift.
- If you have to page or call someone for support, please do so from a
counting house phone or it will be ignored!
- Keep a close eye on the lumi widths as we take data.
- Please continue to take runs even if there is no beam. These data
can be used to determine pedestals. Please label these runs as having no
beam.
- The dithering looks non-invertible. Someone should look at this.
- The Compton polarimeter.
- There are lots of empty shifts that need to be filled immediately!
Please sign up!
Coulombs So Far
- Coulombs taken in this time period: 4.7 C
- Coulombs taken in all so far: 37.5 C as of 7:36am yesterday
Shift Summaries and BTA Reports